Could someone please help me with this? It's my final question on my final exam. I am somewhat stumped.
20th root outside the radical, inside is (-6)^20
Thanks so very much!!

Answer by ccs2011(207)

A negative number raised to an even exponent must be positive
Ex: (-6)^2 = 36
Think of it as 20 negative signs, well thats 10 pairs that cancel themselves out.
Anyways, (-6)^20 = 6^20
Ok now because its a 20th root, rewrite expression using rational exponents.
root is always the denominator, power is the numerator
=> 6^(20/20) = 6^1 = 6
Therefore the solution is 6